| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  84  
 | 
Mon Mar 25 08:24:14 143 | 
Ken Ueberholz | Detectors | XUV-Detector Rates | Count rates were measured for the XUV-
 
detector for a full circle:
 
Cathode: -5V, shields: 0V, main coil 8A, | 
      | 
| 
  237  
 | 
Wed Apr 20 10:09:50 2022 | 
Rodolfo | Laser | Optic with spots | The average energy produced by the laser along 
the beamtime was about 10 mJ/puls. After 
about two weeks of beamtime some of the optics, | 
     | 
| 
  236  
 | 
Sat Apr 16 04:19:05 2022 | 
Patrick, Simon, Alex, Sarper | Runs | Runs night + morning shifts  16.04. | Runs of the night shift 16.04. after changing 
the wavelength range. The new range is 320.12 
- 319.859 nm. The step size was decreased | 
   | 
| 
  235  
 | 
Fri Apr 15 20:43:18 2022 | 
Konstantin | Laser | Dye changed | Dye of the laser system has been changed. 
UV laser power increased from ~8 mJ to 11 
mJ pulse energy. | 
   | 
| 
  234  
 | 
Fri Apr 15 12:53:25 2022 | 
Max, Danyal | Laser | checked laser pulse energy after change in wavelength range | previous scan range: 
 
the laser pulse energy @ 320.62 nm  was ~11 
mJ 
 | 
   | 
| 
  233  
 | 
Fri Apr 15 12:42:18 2022 | 
Max, Danyal, Carsten, Konstantin, Patrick, Simon | General | laser wavelength range changed: 320.12 - 319.60 nm | 
 
Runs of the morning/afternoon/night shift. 
The wavelength range is now 320.12 - 319.60 | 
    | 
| 
  232  
 | 
Thu Apr 14 22:44:40 2022 | 
Patrick, Simon, Danyal, Max | Runs | Runs night shift 14./15.04. | Runs of the night shift. The wavelength range 
is 320.62 - 320.10 nm.
 
 | 
   | 
| 
  231  
 | 
Thu Apr 14 19:46:46 2022 | 
Sarper & Konstantin & Rodolfo | Laser | Laser power | We checked the laser power before starting 
the new scan range.
 
 | 
   | 
| 
  230  
 | 
Thu Apr 14 17:47:55 2022 | 
Carsten, Rodolfo, Konstantin, Sarper | Runs | Afternoon shift runs 14.04 changing the wavelength scan range | The new scan range is changed from [B]321.12 
- 320.60 nm to 320.62 - 320.10 nm[/B]
 
 | 
   | 
| 
  229  
 | 
Thu Apr 14 12:38:38 2022 | 
Sebastian | Detectors | XUV / Cu checkups | Check regularly if the connection to the XUV 
and Cu detectors are still active and the 
detectors move in (after injection) and out | 
   | 
| 
  228  
 | 
Thu Apr 14 12:27:18 2022 | 
Sebastian, Max | Detectors | Connection to XUV / Cu reestablished! | The connection to the XUV and Cu were lost 
during the nightshift.
 
 | 
   | 
| 
  227  
 | 
Thu Apr 14 10:57:11 2022 | 
Carsten | Accelerator | Maintanence UNILAC / Op training | We had no beam since about 1000. Also, we 
are still trying to find out why we have 
such a low intensity in the expereimnt. 
 | 
   | 
| 
  226  
 | 
Thu Apr 14 08:41:42 2022 | 
Carsten | Accelerator | Low intensity at the experiment after series of failures during the night | After the recovery from the failures during 
the night, we see much lower intensity at 
the experiment (measured with the particle | 
   | 
| 
  225  
 | 
Thu Apr 14 06:29:18 2022 | 
Max, Sebastian | Accelerator | lesser ESR injection intensity | We now have less ESR injection intensity (see 
screenshot attached).
 
 | 
    | 
| 
  224  
 | 
Wed Apr 13 15:41:29 2022 | 
Rodolfo & Sarper | Laser | Laser status | We have checked the power of the laser and 
optimized it "a bit".
 
 | 
   | 
| 
  223  
 | 
Wed Apr 13 14:29:04 2022 | 
Carsten, Rodolfo, Sarper, David, Pierre-Michel, Max, Sebastian | Runs | Afternoon - morning shift runs 13.+14.04. | Changed the wavelength scan range from [B]321.62 
- 321.10 nm[/B] to [B][B]321.12 - 320.60[/B] 
nm[/B] after one scan.
 | 
       | 
| 
  222  
 | 
Tue Apr 12 21:56:54 2022 | 
Volker, David, Pierre-Michel & Danyal, Alexandre, Sebastian | General | new laser wavelength scan range 321.62 - 321.10 nm | New wavelength scan range is [B]321.62 - 321.10 
nm[/B]. Before the new scan range (compare 
old scan range in [URL=https://elog.gsi.de/esr/E142/215] | 
     | 
| 
  221  
 | 
Tue Apr 12 17:01:30 2022 | 
Konstantin | Detectors | Voltage PMT North | Voltage of PMT North has been increased from 
2500 V to 2600 V. Additionally the threshold 
was adapted from 25 mV to 12 mV. | 
   | 
| 
  220  
 | 
Tue Apr 12 15:05:56 2022 | 
Rodolfo | Laser | Laser status | We have checked the laser power at 14:15 *
 
 
- Pump energy after Q-Ray @ 532nm: 559 mJ.
 | 
   | 
| 
  219  
 | 
Tue Apr 12 14:52:19 2022 | 
Rodolfo | Detectors | PMT Nord, Middle. Datasheets | Attached are the datasheets for the solar-blind 
PMTs.
 
 | 
    |